A loan is negotiated with the lender agreeing to accept $8, 000 after 1 year, $9, 000 after 2 years, and $20, 000 after four years in full repayment of the loan. The loan is renegotiated so that the borrower makes a single payment of $37, 000 at time T and this results in the same total present value of payments when calculated using an annual effective rate of 7.13%. Find T.
Answer #1

Present value of !st loan at rate = 7.13% annually

PV = 8000/1.0713 + 9000/1.0713^2 + 20000/1.0713^4 = $30493.42

if instead a single payment of $37, 000 at time T, PV of this loan =

PV = 37000/1.0713^T

or same PV of both the loan structure,

37000/1.0713^T = 30493.42

So, T = 2.81 years